A 20-year maturity, 6.5% coupon bond paying coupons semiannually is callable in five years at a call price of $1,010. The bond currently sells at a yield to maturity of 6% (3% per half-year).

a. What is the yield to call annually? (Do not round intermediate calculations. Round your answer to 3 decimal places.)

b. What is the yield to call annually if the call price is only $960? (Do not round intermediate calculations. Round your answer to 3 decimal places.)

c. What is the yield to call annually if the call price is $1,010, but the bond can be called in two years instead of five years? (Do not round intermediate calculations. Round your answer to 3 decimal places.)
Answer #1

fFirst, we need to find the bond's price, for that we need to put the following values in the financial calculator:

INPUT 20*2=40 3 (6.5%/2)*1,000=32.50 1,000
TVM N I/Y PV PMT FV
OUTPUT -1,057.79

Hence, Bond's Price = $1,057.79

a). To find the YTC, we need to put the following values in the financial calculator:

INPUT 5*2=10 -1,057.79 (6.5%/2)*1,000=32.50 1,010
TVM N I/Y PV PMT FV
OUTPUT 2.67

Hence, YTC = 2r = 2 x 2.67% = 5.345%

b). To find the YTC, we need to put the following values in the financial calculator:

INPUT 5*2=10 -1,057.79 (6.5%/2)*1,000=32.50 960
TVM N I/Y PV PMT FV
OUTPUT 2.24

Hence, YTC = 2r = 2 x 2.24% = 4.475%

c). To find the YTC, we need to put the following values in the financial calculator:

INPUT 2*2=4 -1,057.79 (6.5%/2)*1,000=32.50 1,010
TVM N I/Y PV PMT FV
OUTPUT 1.98

Hence, YTC = 2r = 2 x 1.98% = 3.952%